OpenStudy (anonymous):

2w^2-11w+12=0 quadrac formula

OpenStudy (anonymous):

\[2w^{2} - 11w + 12 = 0\]

OpenStudy (anonymous):

\[2w ^{2}-8w-3w+12 =0\] \[2w(w-4)-3(w-4) =0\] \[(2w-3)(w-4) =0\] can you proceed?